OpenStudy (anonymous):

Solve in interval of [0,2pi) 2cosx-3tanx=0

OpenStudy (kirbykirby):

I think you can convert \(\tan x\) into \(\dfrac{\sin x}{\cos x}\), then put everything on the same denominator of \(\cos x\). Then you should be able to convert \(\cos^2 x\) by using \(\cos^2x + \sin^2x=1\) , so that your whole function is in \(\sin x\)'s , and then solve the equation accordingly.

OpenStudy (kirbykirby):

I can provide more guidance if you need it. I guess my help is a bit "vague"

OpenStudy (anonymous):

Yes please. When I put everything under one denominator I have (2cos^2x-3sinx)/cosx. I don't know where to go from there.

OpenStudy (kirbykirby):

Yeah that's good. Try converting \(\cos^2 x\) by using \(1-\sin^2x\) , which comes from \(\cos^2x+\sin^2 = 1\)

OpenStudy (kirbykirby):

\(\sin^2x\) **

OpenStudy (anonymous):

I have (2-sin^2x-3sinx)/cosx, which arranged it would be (-sin^2x-3sinx+2)/cosx

OpenStudy (kirbykirby):

I think you forgot a coefficient of 2 infront of \(\sin^2x\)

OpenStudy (anonymous):

Right.

OpenStudy (kirbykirby):

But once you do that, notice that: \[\frac{-2\sin^2x - 3\sin x +2}{\cos x} =0 \] implies \( -2\sin^2x - 3\sin x +2=0\), And then you say let \(\sin x = y\), and realize that you have a quadratic function that you can solve for \(y\)

OpenStudy (anonymous):

Yes, but what about the cosx?

OpenStudy (kirbykirby):

You can think that you're cross-multiplying your fractions. Like you can say the right side is \(\dfrac{0}{1}\), and you cross multiply so that you get \(1(-2\sin^2 x - 3 \sin x + 2)\) = \(0 *\cos(x)\) , so cos(x) disappears

OpenStudy (anonymous):

Ah, thanks.

OpenStudy (anonymous):

so the solutions I have are pi/6 and 5pi/6

OpenStudy (kirbykirby):

Yeah I get that as well :)

OpenStudy (anonymous):

Awesome, thank you!

OpenStudy (kirbykirby):

=]